OpenStudy (anonymous):

If Logx (1 / 8) = - 3 / 2, then x is equal to

OpenStudy (lgbasallote):

is that read is \[\huge \log_x \left(\frac 18 \right) = -\frac 32\]

OpenStudy (lgbasallote):

is that the question?

OpenStudy (anonymous):

\[\log_{x}8^{-1} =-3/2\] \[-\log_{x}8=-3/2 \]

OpenStudy (anonymous):

\[\log_{x}8=3/2\] \[x^{3/2}=8\] \[x^{3/2}=2^3\] \[x^{1/2}=2\] \[x=\sqrt{2}\]